Can u solve this equation and check results 0.4(2x-3)=1.2(4-x)

Answer by amalm06(224)

0.4(2x-3)=1.2(4-x)
0.8x-1.2=4.8-1.2x
2x=6
x=3 (Answer)
Check: 0.4(6-3)=1.2=1.2(4-3)